A project has the following cash flows:

Year Cash Flows
0 $-16,200
1 $6,900
2 $8,200
3 $6,700

a) What is the NPV at a discount rate of zero percent?

b) What is the NPV at a discount rate of 10 percent?

c) What is the NPV at a discount rate of 20 percent?

d) What is the NPV at a discount rate of 30 percent?
Answer #1

Solution:

NPV = Present value of cash inflows - Present value of cash outflow

a) Discount rate = 0%

NPV = 6900/(1.0)^1 + 8200/(1.0)^2 +6700/(1.0)^3 - 16200= =$5600

b) Discount rate = 10%

NPV = 6900/(1.1)^1 + 8200/(1.1)^2 + 6700/(1.1)^3 -16200*1= $1883.40

c) Discount rate = 20%

NPV = 6900/(1.2)^1 + 8200/(1.2)^2 + 6700/(1.2)^3 - 16200*1= - $878.24

d) Discount rate = 30%

NPV = 6900/(1.3)^1 + 8200/(1.3)^2 + 6700/(1.3)^3 - 16200*1= -$2990.62